1 liter of a gaseous hydrocarbon weighs as much as one liter of co2. Then the molecular formula of hydrocarbon is-

Grade:11

1 Answers

Rahul chouhan
6 years ago
Since weight of hydrocarbon = weight of CO2​Hence, weight of hydrocarbon = 44g/molBase formula = CnHn+r/13Where r = remainderUsing rule 13: Molecular Mass13 = 4413 = C3H3+5 = C3H8Hence molecular formula of compound = C3H8